# Youngjoon Lee1, Jinu Gong2, Joonhyuk Kang1 1School of Electrical Engineering, KAIST, South Korea

### v2

@@ Abstract
Given sufficient data from multiple edge devices, federated learning (FL) enables training a shared model without transmitting private data to a central server. However, FL is generally vulnerable to Byzantine attacks from compromised edge devices, which can significantly degrade the model performance. In this paper, we propose a intuitive plugin that can be integrated into existing FL techniques to achieve Byzantine-Resilience. Key idea is to generate virtual data samples and evaluate model consistency scores across local updates to effectively filter out compromised edge devices. By utilizing this scoring mechanism before the aggregation phase, the proposed plugin enables existing FL techniques to become robust against Byzantine attacks while maintaining their original benefits. Numerical results on medical image classification task validate that plugging the proposed approach into representative FL algorithms, effectively achieves Byzantine resilience. Furthermore, the proposed plugin maintains the original convergence properties of the base FL algorithms when no Byzantine attacks are present.
@@ I Introduction
Recent advances in deep learning have transformed healthcare by leveraging large-scale medical datasets [1, 2]. However, strict patient privacy regulations and HIPAA compliance requirements have created significant barriers to traditional centralized analysis of healthcare data [3]. Federated learning (FL) [4] has emerged as a promising framework for healthcare applications [5]. This approach enables collaborative training of medical AI models while preserving patient privacy through secure parameter sharing [6]. Clinical diagnosis and medical imaging analysis have shown significant improvements using this privacy-preserving paradigm [7]. These healthcare applications handle sensitive protected health information (PHI) that cannot be directly shared among medical institutions [8].
@@
While FL offers privacy benefits, implementing it in real-world healthcare systems faces significant challenges [9]. Healthcare data collected from different medical institutions naturally exhibits data heterogeneity due to varying patient populations and clinical protocols [10, 11]. Moreover, security threats pose a critical concern in federated medical systems, as malicious participants can compromise the global model through Byzantine attacks with poisoned updates [12]. This vulnerability is particularly concerning in clinical applications where model failures could directly impact patient safety and care outcomes [13]. Therefore, developing robust defense mechanisms against adversarial attacks while maintaining model performance under heterogeneous conditions remains an important challenge.
Novel plugin-based architecture proposed as shown in Fig. 1, seamlessly integrates Byzantine resilience into existing FL techniques without compromising their original benefits. Our approach generates virtual data samples to evaluate model behavior patterns. Using these patterns, we compute consistency scores across local updates to identify and filter out malicious edge devices. Our modular approach maintains the core functionality of existing FL techniques. This intuitive plugin can be readily applied into various FL frameworks without requiring complete system redesign. Extensive experiments on medical imaging dataset show that our plugin improves robustness while maintaining high performance across FL algorithms. Main contributions of this paper are as follows:
  • We propose a novel plugin to embed Byzantine resilience into FL techniques without altering their core algorithms.
  • We propose a virtual data-driven scoring method to detect and filter compromised edge devices.
@@ II Preliminaries
In this section, we present background on heterogeneity-aware FL algorithms and their approaches to handling data heterogeneity. To describe FL techniques, we consider a federated network with $K$ edge devices where each device $k$ holds its local dataset $\mathcal{D}_{k}$, FL aims to solve:
@@
where $F_{k}(w)=\frac{1}{|\mathcal{D}_{k}|}\sum_{(x,y)\in\mathcal{D}_{k}}\ell(w;x,y)$ represents the local objective function of device $k$.
@@
where $g_{e}$ denotes global epoch and $\mu$ controls client drift. This proximal regularization allows FedProx to achieve more stable convergence even when data across devices is highly heterogeneous.
FedDyn [15] utilizes dynamic regularization to align local optimization objectives with the global goal by adapting to local optimization paths. The local objective in FedDyn are as follows:
$F_{k}^{Dyn}(w)=F_{k}(w)+h^{g_{e}}(w^{g_{e}})^{T}(w-w^{g_{e}})+\frac{\alpha}{2}|w-w^{g_{e}}|^{2},$ (3)
where $\alpha$ is a scaling factor and $h^{g_{e}}$ captures optimization trajectory differences. This adaptive approach allows the model to dynamically update regularization, ensuring a better fit to each device’s unique data distribution.
@@
FedSAM [17] incorporates Sharpness-Aware Minimization (SAM) to enhance model generalization on heterogeneous data. SAM modifies the objective to minimize sharp local optima by perturbing gradients:
$w_{k}^{SAM}=w_{k}^{l_{e}}-\rho\frac{\nabla F_{k}(w_{k}^{l_{e}}+\epsilon)}{|\nabla F_{k}(w_{k}^{l_{e}}+\epsilon)|},$ (5)
where $l_{e}$ denotes local epoch and $\epsilon=\rho\frac{\nabla F_{k}(w_{k}^{l_{e}})}{|\nabla F_{k}(w_{k}^{l_{e}})|}$ is a perturbation that stabilizes convergence by directing models toward flatter minima.
FedSpeed [18] accelerates convergence and enhances consistency through prox-correction and gradient perturbation:
@@ III-A Byzantine Attack and Non-IID Setting
In this paper, we consider FL environment vulnerable to both targeted and untargeted Byzantine attacks in medical imaging task as [19]. Specifically, the healthcare FL system comprises $K$ participating edge devices, including $B$ benign and $M$ compromised nodes, all connected to a central server. Each edge device $k$ maintains its private patient dataset $\mathcal{D}_{k}$ with varying sizes. To preserve patient privacy while developing a shared medical image classification model, edge devices collaborate through $G$ global epochs by sharing only model parameters with the central server.
@@ III-B Plugin-based Byzantine-Resilient FL
In this section, we introduce a plugin-based approach that enhances FedAvg’s resilience against Byzantine attacks through feature-space analysis. At first, benign edge devices execute local learning and malicious edge devices craft byzantine attacks. For benign edge devices $b\in B$, local training aims to minimize the empirical loss:
@@
where $\ell(\cdot)$ is the cross-entropy loss function. The local model updates are performed via SGD with learning rate $\eta$:
$w_{b}^{l_{e}+1}=w_{b}^{l_{e}}-\eta\nabla F_{b}(w_{b}^{l_{e}};\mathcal{B}_{b}),$ (8)
@@
Figure 2: Illustration of the proposed plugin’s scoring mechanism in feature space. The plugin measures pairwise cosine similarities (denoted as ‘cos’) between feature vectors extracted from model updates ($f_{1}$ through $f_{5}$) relative to the global model ($f_{g}$). Benign models (blue) exhibit high cosine similarity scores with other benign models, while poisoned models (red) show distinct patterns, enabling effective filtering of Byzantine attacks.
However, compromised edge devices $m\in M$ may perform malicious updates through either targeted or untargeted attacks. In targeted attacks, compromised devices modify their local updates which finished $L$ local training, to influence specific samples while maintaining overall performance:
@@
where $\lambda$ is a boosting factor designed to amplify the attack’s impact. For untargeted attacks, malicious devices inject arbitrary noise to degrade overall model performance:
@@
To defend against these attacks, our plugin performs Byzantine attack filtering using feature space analysis in three main steps: deviation analysis, feature mapping, and similarity-based rejection. First, we generate $N$ virtual samples $\{v_{n}\}_{n=1}^{N}$ from a standard normal distribution $\mathcal{N}(\bm{0},I)$ to serve as probe points for analyzing model behaviors. For each model pair $(w_{i}^{g_{e}},w_{j}^{g_{e}})$, we compute their deviations from the global model $w^{g_{e}}$:
$\Delta w_{i}$ $=w_{i}^{g_{e}}-w^{g_{e}}$ (13)
$\Delta w_{j}$ $=w_{j}^{g_{e}}-w^{g_{e}}$ (14)
@@
Next, we map these deviations to feature representations using a feature extractor $\mathbf{g}_{\phi}(\cdot)$, which is implemented as all layers of the base model except the final classification layer. For each model pair $(w_{i},w_{j})$ and virtual sample $v_{n}$, we compute:
$f_{i}$ $={f_{i}^{1},f_{i}^{2},...,f_{i}^{N}}=\mathbf{g}_{1:L-1}(v_{n};\Delta w_{i})$ (15)
$f_{j}$ $={f_{j}^{1},f_{j}^{2},...,f_{j}^{N}}=\mathbf{g}_{1:L-1}(v_{n};\Delta w_{j})$ (16)
where $\mathbf{g}_{1:L-1}$ denotes the feature extraction layers of the model (all layers except the final classification layer), and each $f_{i}^{n},f_{j}^{n}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature vector for the $n$-th virtual sample.
After that, pairwise similarity between models is then quantified using the average cosine similarity across all virtual samples:
$s_{i,j}=\frac{1}{N}\sum_{n=1}^{N}\cos(f_{i}^{n},f_{j}^{n})$ (17)
where $\cos(f_{i}^{n},f_{j}^{n})$ measures the angular similarity between feature vectors, with higher values indicating more similar behavior patterns.
As shown in Fig. 2, benign models naturally cluster together with high similarity scores, while Byzantine models exhibit distinct patterns. For each model update $w_{k}$, we calculate its average cosine similarity with other updates:
$\bar{s}_{k}=\frac{1}{K-1}\sum_{j\neq k}s_{k,j}$ (18)
We then sort these average similarity scores $\{\bar{s}_{k}\}_{k=1}^{K}$ in ascending order and define set $\mathcal{S}$ as:
$\mathcal{S}=\{k\mid\pi(\bar{s}_{k})>M\}$ (19)
where $\pi(\bar{s}_{k})$ represents the position of $\bar{s}_{k}$ in the sorted scores, effectively excluding the $M$ updates with lowest similarity scores as potential Byzantine attacks. Subsequently, Byzantine-resilient global model is then updated through averaging of the remaining model updates:
$w^{g_{e}+1}=\frac{1}{|\mathcal{S}|}\sum_{k\in\mathcal{S}}w_{k}$ (20)
Lastly, central server broadcast the aggregated model to all edge devices for the next $g_{e}$ of local training. This process continues iteratively until the desired model performance is achieved or a maximum number of $g_{e}$ is reached. Overall procedure of proposed plugin is described in Algorithm 1.
@@
/* Virtual data-drive consistency scoring at central server */
@@ 1 Generate $\{v_{n}\}_{n=1}^{N}\sim\mathcal{N}(\bm{0},I)$;
4       $f_{i}\leftarrow\mathbf{g}_{1:L-1}(v_{n};w_{i}^{g_{e}}-w^{g_{e}})$;
5       $f_{j}\leftarrow\mathbf{g}_{1:L-1}(v_{n};w_{j}^{g_{e}}-w^{g_{e}})$;
@@ 11 $\mathcal{S}=\{k\mid\pi(\bar{s}_{k})>M\}$;
Algorithm 1 Proposed Virtual Data-driven Byzantine Attack Filtering Plugin
@@ IV-A Experiment setting
We evaluate our proposed plugin on blood cell classification task [20], implementing a ResNet-18 [21] architecture as our base model. Additionally, Gaudi 2 AI accelerator from Intel was deployed to efficiently manage the computationally demanding process, delivering scalable performance throughout our experimental work. Our experiments compare the performance of various FL techniques (FedAvg, FedProx, FedDyn, FedRS, FedSAM, FedSpeed) with and without our proposed plugin under both Targeted Model Poisoning (TMP) and Untargeted Model Poisoning (UMP) attack scenarios. To simulate realistic data heterogeneity, we distribute the data non-uniformly across $K=10$ medical edge devices following the quantity skew protocol [22]. The complete implementation details and configuration parameters are publicly available in our repository [^1] https://github.com/NAVER-INTEL-Co-Lab/gaudi-byzantine.
@@ IV-B1 Impact of Byzantine Attacks
First, we check the effect of model poisoning attacks on representative heterogeneity-aware FL techniques. Experimental results reveal the significant vulnerability of FL methods to Byzantine attacks as shown in Fig. 3. Despite handling data heterogeneity effectively, methods like FedProx, FedRS, FedDyn, FedSAM, and FedSpeed show severe performance degradation. These methods exhibit vulnerability under both targeted and untargeted attacks in our experiments. As the ratio of compromised devices increases, we observe a dramatic decrease in accuracy across all methods. Performance drops to approximately 20% when $p=0.3$, indicating their inherent weakness against malicious updates without proper defense mechanisms.
Furthermore, performance degradation varies between attack types. Untargeted attacks show notably more severe impact compared to targeted attacks across all FL methods. These methods maintain acceptable accuracy under targeted attacks with lower compromise ratios ($p=0.1$). However, they suffer catastrophic failure under untargeted attacks even at these lower ratios. This vulnerability originates from their primary focus on addressing data heterogeneity. As a result, they lack effective defenses against malicious updates in practical FL systems.
Additionally, we check conventional Byzantine-Resilient methods like Krum [23], Trimmed-Mean [24], and Fang [25] as baselines. These methods help demonstrate the relative impact of Byzantine attacks on heterogeneity-aware FL methods. The performance gap between these two categories is significant. This difference clearly emphasizes the severity of Byzantine attacks in heterogeneous FL environments.
(a) targeted attack
(b) untargeted attack
Figure 3: Performance comparison of heterogeneity-aware FL methods and Byzantine-Resilient FL methods under different attack scenarios. The x-axis represents the ratio of compromised devices ($p$), and the y-axis shows the maximum test accuracy (%).

### v3

@@ Abstract
Given sufficient data from multiple edge devices, federated learning (FL) enables training a shared model without transmitting private data to the central server. However, FL is generally vulnerable to Byzantine attacks from compromised edge devices, which can significantly degrade the model performance. In this work, we propose an intuitive plugin that seamlessly embeds Byzantine resilience into existing FL methods. The key idea is to generate virtual data samples and evaluate model consistency scores across local updates to effectively filter out compromised updates. By utilizing this scoring mechanism before the aggregation phase, the proposed plugin enables existing FL methods to become robust against Byzantine attacks while maintaining their original benefits. Numerical results on blood cell classification task demonstrate that the proposed plugin provides strong Byzantine resilience. In detail, plugin-attached FedAvg achieves over 89.6% test accuracy under 30% targeted attacks (vs. 19.5% w/o plugin) and maintains 65–70% test accuracy under untargeted attacks (vs. 17–19% w/o plugin).
@@ I Introduction
Recent advances in deep learning have transformed healthcare by leveraging large-scale medical datasets [1, 2]. However, strict patient privacy regulations and HIPAA compliance requirements have created significant barriers to traditional centralized analysis of healthcare data [3]. Federated learning (FL) [4] has emerged as a promising framework for healthcare applications [5]. This method enables collaborative learning of medical AI models while preserving patient privacy through secure parameter sharing [6]. For example, clinical diagnosis and medical imaging analysis have shown significant improvements using this privacy-preserving paradigm [7].
@@
While FL offers privacy benefits, implementing it in real-world healthcare systems faces significant challenges [8]. Healthcare data collected from different medical institutions naturally exhibits data heterogeneity due to varying patient populations and clinical protocols [9, 10]. Moreover, Byzantine threats pose a critical concern in federated medical systems, as malicious participants can compromise the global model through Byzantine attacks with poisoned updates [11]. This vulnerability is particularly concerning in clinical applications where model failures could directly impact patient safety and care outcomes [12]. Therefore, developing Byzantine robust FL against adversarial attacks while maintaining model performance under heterogeneous conditions remains an important challenge.
The novel plugin-based architecture proposed in Fig. 1 embeds Byzantine resilience into existing FL methods without compromising their original benefits. Our plugin generates virtual data samples to evaluate model behavior patterns, computing consistency scores across local updates to detect and filter malicious contributions while preserving the core functionality of FL methods. Extensive experiments on medical imaging datasets demonstrate that this approach enhances robustness while maintaining high performance across diverse FL methods. Main contributions of this paper are as follows:
  • We propose a novel plugin to embed Byzantine resilience into FL methods without altering their core principle.
  • We propose a virtual data-driven scoring method to detect and filter compromised local updates.
@@
The remainder of this paper is organized as follows. Section II provides background on representative FL methods. Section III presents our novel modular plugin for FL. Section IV demonstrate how our plugin enhances performance against Byzantine attacks through extensive numerical evaluations.
@@ II Preliminaries
To describe FL methods, we consider a federated network with $K$ edge devices, where each device $k$ holds a local dataset $\mathcal{D}_{k}$. The goal of FL is to solve:
@@
where $F_{k}(w)=\frac{1}{|\mathcal{D}_{k}|}\sum_{(x,y)\in\mathcal{D}_{k}}\ell(w;x,y)$ represents the local objective function of edge device $k$.
@@
where $g_{e}$ denotes global epoch and $\mu$ controls the client drift. This proximal regularization allows FedProx to achieve more stable convergence even when data across devices is highly heterogeneous.
FedDyn [14] utilizes dynamic regularization to align local optimization objectives with the global goal by adapting to local optimization paths. The local objective in FedDyn is as follows:
$F_{k}^{Dyn}(w)=F_{k}(w)+(h^{g_{e}})^{T}(w-w^{g_{e}})+\frac{\alpha}{2}|w-w^{g_{e}}|^{2},$ (3)
where $\alpha$ is a scaling factor and $h^{g_{e}}$ captures optimization trajectory differences. Then, the central server aggregates local updates and adjusts the global model using dynamic control variates. This adaptive approach allows the model to dynamically update regularization, ensuring a better fit to each device’s unique data distribution.
@@
FedSAM [16] incorporates Sharpness-Aware Minimization (SAM) [17] to enhance model generalization on heterogeneous data. SAM modifies the objective to minimize sharp local optima by perturbing gradients:
$w_{k}^{SAM}=w_{k}^{g_{e},l_{e}}-\rho\frac{\nabla F_{k}(w_{k}^{g_{e},l_{e}}+\epsilon)}{|\nabla F_{k}(w_{k}^{g_{e},l_{e}}+\epsilon)|},$ (5)
where $l_{e}$ denotes local epoch and $\epsilon=\rho\frac{\nabla F_{k}(w_{k}^{g_{e},l_{e}})}{|\nabla F_{k}(w_{k}^{g_{e},l_{e}})|}$ is a perturbation that stabilizes convergence by directing models toward flatter minima.
FedSpeed [18] accelerates convergence through prox-correction and gradient perturbation:
@@ III-A Byzantine Attack and Non-IID Setting
We consider an FL environment vulnerable to both targeted and untargeted model poisoning attacks in medical imaging task as [19]. Specifically, the FL system comprises $K$ participating edge devices, including $B$ benign and $M$ compromised nodes, all connected to a central server. Each edge device $k$ maintains its private patient dataset $\mathcal{D}_{k}$ with varying sizes. To preserve patient privacy, edge devices collaboratively train a shared medical image classifier over $G$ global epochs by sharing only model parameters with the central server.
@@ III-B Plugin-based Byzantine-Resilient FL
In this section, we introduce a plugin-based approach that enhances FedAvg’s resilience against Byzantine attacks through feature-space analysis. First, benign edge devices perform local learning, while malicious edge devices craft Byzantine attacks. For benign edge devices $b\in B$, local training aims to minimize the empirical loss:
@@
where $\ell(\cdot)$ is the cross-entropy loss function. The local updates are performed via SGD with learning rate $\eta$:
$w_{b}^{g_{e},l_{e}+1}=w_{b}^{g_{e},l_{e}}-\eta\nabla F_{b}(w_{b}^{g_{e},l_{e}};\mathcal{B}_{b}),$ (8)
@@
However, compromised edge devices $m\in M$ may perform malicious updates through either targeted or untargeted attacks. In targeted attacks, compromised devices alter their local updates after $L$ local training epochs, to mislead specific samples while preserving overall performance.
@@
where $\lambda$ is a boosting factor designed to amplify the attack’s impact. For untargeted attacks, malicious devices inject arbitrary noise to degrade overall performance:
@@
Figure 2: Illustration of the proposed plugin’s scoring principle. The plugin measures pairwise cosine similarities (denoted as ‘cos’) between feature vectors of model updates ($f_{1}$ through $f_{5}$) relative to the global model ($f_{g}$). Benign models (blue) have high cosine similarity with each other, whereas poisoned models (red) show distinct patterns, enabling effective Byzantine attack filtering.
To filter these attacks, our plugin performs Byzantine attack filtering using feature space analysis in three main steps: deviation analysis, feature mapping, and similarity-based rejection. First, the central server generate $N$ virtual samples $\{v_{n}\}_{n=1}^{N}$ from a standard normal distribution $\mathcal{N}(\bm{0},I)$ to serve as probe points for analyzing model behaviors. For each model pair $(w_{i}^{g_{e}},w_{j}^{g_{e}})$, the server compute their deviations from the global model as:
$\Delta w_{i}=w_{i}^{g_{e}}-w^{g_{e}},\;\Delta w_{j}=w_{j}^{g_{e}}-w^{g_{e}}.$ (13)
@@
Next, the central server maps these deviations to feature representations using a feature extractor $g_{\phi}(\cdot)$, which is implemented as all layers of the base model except the final classification layer. For each model pair $(w_{i},w_{j})$ and virtual sample $v_{n}$, the server computes:
$\mathcal{F}_{i}$ $=\{f_{i}^{1},f_{i}^{2},\dots,f_{i}^{N}\}=g_{1:L-1}(v_{n};\Delta w_{i}),$ (14)
$\mathcal{F}_{j}$ $=\{f_{j}^{1},f_{j}^{2},\dots,f_{j}^{N}\}=g_{1:L-1}(v_{n};\Delta w_{j}),$ (15)
where $g_{1:L-1}$ denotes the feature extraction layers of the model (all layers except the final classification layer), and each $f_{i}^{n},f_{j}^{n}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature vector for the $n$-th virtual sample.
The server then computes the pairwise similarity between models using the average cosine similarity over the $N$ virtual samples:
$s_{i,j}=\frac{1}{N}\sum_{n=1}^{N}\cos\big{(}f_{i}^{n},f_{j}^{n}\big{)},\ f_{i}^{n}\in\mathcal{F}_{i},~f_{j}^{n}\in\mathcal{F}_{j}.$ (16)
where $\cos(f_{i}^{n},f_{j}^{n})$ measures the angular similarity between feature vectors, with higher values indicating more similar behavior patterns. As shown in Fig. 2, benign models naturally cluster together with high similarity scores, whereas malicious model updates exhibit distinct patterns. For each model update $w_{k}$, the server calculates its average cosine similarity with the other updates as:
$\bar{s}_{k}=\frac{1}{K-1}\sum_{j\neq k}s_{k,j}.$ (17)
The server then sorts these average similarity scores $\{\bar{s}_{k}\}_{k=1}^{K}$ in ascending order and defines the set $\mathcal{S}$ as:
$\mathcal{S}=\{k\mid\pi(\bar{s}_{k})>M\},$ (18)
where $\pi(\bar{s}_{k})$ denotes the position of $\bar{s}_{k}$ in the sorted scores, effectively excluding the $M$ model updates with the lowest similarity scores as potential Byzantine attacks. The Byzantine-resilient global model is then updated by averaging the remaining model updates:
$w^{g_{e}+1}=\frac{1}{|\mathcal{S}|}\sum_{k\in\mathcal{S}}w_{k}.$ (19)
Finally, the central server broadcasts the aggregated model to all edge devices for the next global training round. The overall procedure of the proposed plugin for FL method is described in Algorithm 1.
@@
/* Virtual data-driven consistency scoring at the central server */
@@ 1 Generate $\{v_{n}\}_{n=1}^{N}\sim\mathcal{N}(\bm{0},I)$;
4 $\mathcal{F}_{i}\leftarrow\mathbf{g}_{1:L-1}(v_{n};\Delta w_{i})$;
5 $\mathcal{F}_{j}\leftarrow\mathbf{g}_{1:L-1}(v_{n};\Delta w_{j})$;
@@ 11 $\mathcal{S}=\{k\mid\pi(\bar{s}_{k})>M\}$;
Algorithm 1 Proposed Modular Plugin for FL
@@ IV-A Experiment setting
We evaluate our proposed plugin on blood cell classification task [20], where each edge device runs a ResNet-18 [21] model as its local model. Our experiments compare the performance of various FL methods with and without our proposed plugin under both Targeted Model Poisoning (TMP) and Untargeted Model Poisoning (UMP) attack scenarios. To simulate realistic data heterogeneity, we distribute the data non-uniformly across $K=10$ medical edge devices following the quantity skew protocol [22]. The complete implementation details and configuration parameters are publicly available in our open repository [^1] https://github.com/yjlee22/fl-plugin.
@@ IV-B1 Impact of Byzantine Attacks
To examine the vulnerability of heterogeneity-aware FL methods, we check the performance under model poisoning attacks. As shown in Fig. 3, all vanilla FL methods experience severe degradation under Byzantine attack settings. As the proportion of compromised devices increases, their accuracy falls sharply, reaching about 20% when $p=0.3$. This trend highlights the absence of mechanisms to identify and remove malicious updates. Moreover, untargeted attacks result in greater harm than targeted attacks even when the compromise ratio is low. In detail, all heterogeneity-aware FL methods achieve less than 20% accuracy when subjected to untargeted attacks. Thus, additional Byzantine resilient components are essential for practical FL deployments.
In addition, we compare with representative Byzantine-resilient FL methods, including Krum [23], Trimmed-Mean [24], and Fang [25]. These Byzantine-resilient FL methods consistently outperform heterogeneity-aware ones in adversarial settings. The noticeable performance gap underscores the severity of Byzantine threats in realistic heterogeneous networks. Furthermore, untargeted attacks remain highly disruptive due to their random and unpredictable perturbations. Hence, secure FL requires mechanisms beyond addressing data heterogeneity to counter adaptive attackers.
(a) Targeted Model Poisoning Attack
(b) Untargeted Model Poisoning Attack
Figure 3: Performance comparison of heterogeneity-aware and Byzantine-resilient FL methods under model poisoning attacks. The x-axis denotes the ratio of compromised devices ($p$), and the y-axis shows the test accuracy (%).